# The effect of accretion on scalar superradiant instability

### v1

@@ III.2 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$
Table 1: Comparison of the analytical estimates to the numerical results obtained from solving Eqs. (11) for essential quantities. Initial parameters are identical to those used in Fig. 2. Since the condensate mass does not present a maximum in the attractor phase for ${f}_{\mathrm{Edd}}=0$, $10^{-17}$ and $10^{-15}$, the corresponding quantities at $t_{2}$ are labeled with “N/A” in the table. Moreover, $t_{3}$ of ${f}_{\mathrm{Edd}}=0$ case is not applicable, which is also labeled with “N/A”.
III.3 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{SR}}^{(011)}$
@@
The spin-up phase ends at $t_{0}^{\prime}$, when the BH spin reaches the local maximum for the first time. Expanding the first term on the right side of Eq. (40) at $a_{*}=1$ and combining it with Eq. (41), Eq. (40) reduces to
$\begin{split}\dot{a}_{*}=&\frac{3\cdot 2^{1/3}}{\tau_{\text{acc}}}\left(1-a_{*}\right)^{2/3}\\ &-\frac{M_{\text{s,0}}}{24}M^{6}\mu^{8}\exp\left[\frac{\tau_{\text{acc}}}{192}(M^{8}-M^{8}_{0})\mu^{9}\right].\end{split}$ (57)
@@
$\begin{split}\alpha_{0^{\prime}}=&\left(\frac{144}{\tau_{\mathrm{acc}}\mu}\right)^{{1}/{8}}\\ &\times\left[W_{0}\left(2^{{4}/{9}}3^{{2}/{3}}M_{\text{s,0}}^{-{4}/{3}}\tau_{\text{acc}}^{-{1}/{3}}\mu^{-{5}/{3}}e^{\frac{\alpha_{0}^{8}\tau_{\mathrm{acc}}\mu}{144}}\right)\right]^{{1}/{8}},\end{split}$ (58)
@@
The time $t_{0}^{\prime}$ can then be calculated with Eq. (19), and the BH spin and condensate mass at $t_{0}^{\prime}$ can be derived by Eqs. (13) and (37), respectively. The subsequent evolution is the same as the scenario described in Sec. III.2, with the subscript $0^{\prime\prime}$ in Eq. (44) replaced by $0^{\prime}$.
@@
Actually, Eq. (58) could be used to judge which one of the three groups the evolution should be attributed to. If the obtained $\alpha_{0^{\prime}}$ is larger than $\sqrt{6}k\alpha_{0}$, one should instead use the analysis in Sec. III.2. On the other hand, if $\alpha_{0^{\prime}}$ is smaller than $\alpha_{0}$, the evolution does not follow the accretion trajectory at all, which case is discussed in Sec. III.4 below.
III.4 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\gg{\tau}_{\mathrm{SR}}^{(011)}$
@@
This group includes three different scenarios ${\tau}_{\mathrm{SR}}^{(011)}\ll{\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{GW}}^{(011)}$, ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{GW}}^{(011)}$ and ${\tau}_{\mathrm{acc}}\gg{\tau}_{\mathrm{GW}}^{(011)}$, corresponding to ${f}_{\mathrm{Edd}}=10^{-9}$, $10^{-15}$ and $10^{-17}$ in Fig. 2, respectively. The curves with $\tau_{\text{acc}}\gg\tau_{\text{SR}}^{(011)}$ deviate from the $f_{\text{Edd}}=0$ case only from $t\sim\tau_{\text{acc}}$. The whole evolution can also be split into different phases. Specially, compared to the two cases analysed in Sections III.2 and III.3, the spin-up phase vanishes when $\tau_{\text{acc}}\gg\tau_{\text{SR}}^{(011)}$ and $a_{*0}>a_{*\text{c}}^{(011)}$.
@@
The BH spin drops from $t=0$ and reaches the minimum at time $t_{1}$. The BH mass $M_{1}$ and the condensate mass $M_{\text{s,1}}$ can be calculated in the same way as in Sec. II.2.2. After $t_{1}$, the BH follows the Regge trajectory. The condensate mass may either increase or decrease, depending on the relative size of the superradiance rate and the GW emission rate at $t_{1}$. No matter which case, Eq. (51) can be used to estimate $\alpha_{2}$. If $\alpha_{2}>\alpha_{1}$, the condensate continues to grow after $t_{1}$ and $M_{\mathrm{s}}^{(011)}/M$ reaches its maximum value at $t_{2}$. Otherwise, the condensate shrinks since $t_{1}$. In the case of $\alpha_{2}>\alpha_{1}$, the time $t_{2}$ can be further calculated using Eq. (54), and the condensate mass at $t_{2}$ can be estimated with Eq. (49).
III.5 $a_{*0}<a_{*\text{c}}^{(011)}$
@@
When $a_{*0}$ is smaller than the critical value $a_{\text{c}}^{(011)}$, the superradiance condition is not satisfied initially. Nonetheless, the evolution of the BH-condensate system can still be split into the four phases explained above, as shown in Fig. 4. The angular momentum and energy flow from the condensate to the BH at first. Since the initial condensate mass is much smaller than the BH mass, this contribution is unimportant and the BH mass evolution is dominated by accretion. As a result, the growth of the BH mass can be estimated with Eq. (19). The BH spin also increases, following the relation in Eq. (13). Eventually, the BH spin is above the critical value and the superradiance is turned on. When superradiance balances accretion, the BH spin reaches its local maximum, which is the end of the spin-up phase. After the local maximum of the BH spin, the BH spin drops quickly to the Regge trajectory and then evolves along it afterward. The condensate mass first drops before the BH spin grows to its critical value $a_{*\text{c}}^{(011)}$. It can be estimated using Eq. (37) during this time interval. Since then, the system evolves as in the $a_{*0}>a_{\mathrm{*c}}^{(011)}$ cases, which are explained in previous subsections.
@@
In Fig. 4, we also compared the curves with $a_{*0}=0$ to those with $a_{*0}=0.7$. Because the condensate initial mass is small, the evolutions of BH mass are indistinguishable between these two cases, confirming our argument that the BH mass growth is dominated by accretion. In contrast, the BH spin and condensate mass evolution have different behavior with different values of $a_{*0}$. Especially, the spin-up and spin-down phases are largely affected by the initial BH spin. The corresponding estimates are similar to those in Sec. III.3, with a revision to the calculation of $\alpha_{0^{\prime}}^{(011)}$. When $a_{*0}<a_{\mathrm{*c}}^{(011)}$, the spin-up phase can be divided into two segments, separated by the time $t_{\mathrm{int}}$ when the BH spin reaches $a_{*\text{c}}^{(011)}$. The BH mass at this time can be obtained by solving Eq. (10) together with Eq. (13). The mass coupling at this time, defined as $\alpha_{\text{int}}$, can be accomplished either numerically or analytically in powers of $\alpha_{0}$
$\frac{\alpha_{\text{int}}}{\alpha_{0}}=k+k\sum_{n=1}^{+\infty}c_{n}\left(\frac{k\alpha_{0}}{m}\right)^{n},$ (59)
where
$c_{1}$ $=2^{5/2}3^{-3/2},$ (60a)
$c_{2}$ $=94/27,$ (60b)
$c_{3}$ $=\frac{2^{3/2}}{3^{7/2}}\left(172-9\frac{m^{2}}{\bar{n}^{2}}\right),$ (60c)
$c_{4}$ $=\frac{2}{729}\left(12451-1701\frac{m^{2}}{\bar{n}^{2}}\right).$ (60d)
@@
With the obtained $\alpha_{\text{int}}$, the condensate mass at $t_{\text{int}}$ can be estimated with Eq. (37). Beyond $t_{\text{int}}$, the formulas obtained above for the $a_{*0}>a_{\text{*c}}^{(011)}$ scenario can be applied. One could simply replace the $\alpha_{0}$, $a_{*0}$ and $M_{\text{s,0}}$ by $\alpha_{\text{int}}$, $a_{*\text{c}}^{(011)}$ and $M_{\text{s}}(t_{\text{int}})$, respectively.
@@
It is noteworthy that if the initial condensate mass is sufficiently large, the BH spin reaches ${a}_{\mathrm{*c}}^{(011)}$ at a faster rate with the additional contribution from the scalar field. Although this scenario can be easily adapted from Sec. III.4, it is beyond the scope of this work. If modes with higher $l$ are considered, the BH spin may not reach $a_{*\text{c}}^{(011)}$. This scenario will be discussed in the next section.
@@
III.6 Evolution on the Regge plot
Figure 5: The evolution of the BH spin as a function of the BH mass. The initial parameters for panels (a) and (b) are identical to those used in Fig. 2 and Fig. 4, respectively. The $\{0,1,1\}$ Regge trajectory defined in Eq. (10) is represented by the upper boundary of the shaded region.
Despite the complications in the evolution of the BH-condensate systems with different accretion rates, the evolution presents a rather simple and universal behavior on the Regge plot. In Fig. 5(a), we show the evolutions of the system with three typical values of $f_{\text{Edd}}$, corresponding to the three scenarios detailed in Sec. III.2, Sec. III.3 and Sec. III.4, respectively. The initial BH spin and mass are $a_{*0}=0.7$ and $M_{0}=10^{3}M_{\odot}$, respectively. The initial condensate mass is $M_{\text{s},0}=10^{-5}M_{0}$ and the initial mass coupling is $\alpha_{0}=0.01$. In the evolution, only the $\{0,1,1\}$ mode is considered. The GW emission of the spinning condensate is also included. In Fig. 5(b), the initial BH spin is $a_{*0}=0$, corresponding to the scenario detailed in Sec. III.5.
@@ III.6 Evolution on the Regge plot
We then turn to the case with the fastest accretion, regardless of whether $a_{*0}>a_{*\mathrm{c}}^{(011)}$, corresponding to $f_{\text{Edd}}=1$ in Fig. 5. The accretion dominates the evolution of the BH and the superradiance can be safely ignored initially. The BH first follows the trajectory determined by Eq. (20), referred to as accretion trajectory below. The BH spin approaches $a_{*\text{lim}}$ and the BH mass grows almost exponentially. During this time, the condensate grows in mass and angular momentum. At some time, the condensate is dense enough such that the superradiance surpasses the accretion, leading to the fast drop to the Regge trajectory. The evolution follows the Regge trajectory afterward.
@@
From the Regge plot, it is clear that the entire evolution has a universal behavior. It consists of the parts following the two trajectories and the transition between them. They correspond to the spin-up, spin-down, and attractor phases in Sec. III.2. The decaying phase could not be observed in the BH Regge plot. It is clear by looking at the evolution of the condensate mass.
@@
Due to this universal behavior, one only needs to identify the BH masse $M$ or spin $a_{*}$ at these critical points at which the system leaves the accretion trajectory and merges to the Regge trajectory. To determine these key points and the evolution between them, the following procedure can be applied. If $M_{0}\omega_{011}\geq 1/2$, superradiance does not occur and the system follows the accretion trajectory. If $M_{0}\omega_{011}<1/2$, $\alpha_{0^{\prime}}$ can be calculated using Eq. (58), as in Sec. III and Sec. III.5. Based on the obtained $\alpha_{0^{\prime}}$, there are three possible cases:
@@
  • $\alpha_{0^{\prime}}>\sqrt{6}k\alpha_{0}$: This case corresponds to the scenario in Sec. III.2 and the curves with $f_{\text{Edd}}=1$ in Fig. 5. The system initially follows the accretion trajectory approximately described by Eq. (13). Then it rapidly transits to the Regge trajectory in Eq. (10) at $\alpha_{1}$ given in Eq. (44). After that, the system follows the $\{0,1,1\}$ Regge trajectory until $M=M_{\mathrm{de}}^{(011)}$.
@@
  • $\alpha_{0}<\alpha_{0^{\prime}}<\sqrt{6}k\alpha_{0}$: This case corresponds to the scenario in Sec. III.3 and the curves with $f_{\text{Edd}}=10^{-3}$ in Fig. 5. The system first follows the accretion trajectory and leaves this trajectory at $\alpha_{0^{\prime}}$. It then merges onto the Regge trajectories at $\alpha_{1}$ given in Eq. (44). Then it follows this trajectory until $M=M_{\mathrm{de}}^{(011)}$.
@@
  • $\alpha_{0^{\prime}}<\alpha_{0}$: This case corresponds to the scenario in Sec. III.4 and the curves with $f_{\text{Edd}}=10^{-9}$ in Fig. 5. The system is dominated by superradiance initially and quickly merges to the Regge trajectories at $M=M_{1}$. Here $M_{1}$ is estimated using Eqs. (21) and (23b). The system then follows the $\{0,1,1\}$ Regge trajectory until $M=M_{\mathrm{de}}^{(011)}$.
@@
Finally, we turn to what is not reflected in the Regge plots. The BH mass grows almost exponentially before merging onto the Regge trajectory, from which one could obtain the corresponding time cost. This can then be applied to calculate the condensate mass and GW emission rate. When the BH evolves along the Regge trajectory, the Eqs. (45) to (56) can be directly applied for estimating the important quantities in this phase.
@@
With these equations, one could understand why the condensate mass evolution and GW emission flux are dramatically affected when $\tau_{\text{acc}}\lesssim\tau_{\text{SR}}$, which is shown in Fig. 2. Without accretion, the BH mass $M$ is roughly constant, resulting in an exponential increase of the condensate mass. The time duration of the GW is controlled by $\tau_{\text{GW}}$. Comparably, the condensate mass with accretion is a double exponential function and the termination of the GW emission is controlled by $\tau_{\text{de}}\ll\tau_{\text{GW}}$. The maximum condensate mass is also larger with accretion, which reaches as much as 0.15 in Fig. 3, while this number never exceeds 0.1 if accretion is absent. Consequently, the GW emission in Eq. (82) is stronger in magnitude but shorter in time.
IV Multi-mode evolution with accretion

### v2

@@ 3.2 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$
Table 1: Comparison of the analytical estimates to the numerical results obtained from solving Eqs. (2.11) for essential quantities. Initial parameters are identical to those used in Fig. 2. Since the condensate mass does not present a maximum in the attractor phase for ${f}_{\mathrm{Edd}}=0$, $10^{-17}$ and $10^{-15}$, the corresponding quantities at $t_{2}$ are labeled with “N/A” in the table. Moreover, $t_{3}$ of ${f}_{\mathrm{Edd}}=0$ case is not applicable, which is also labeled with “N/A”.
3.3 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{SR}}^{(011)}$
@@
The spin-up phase ends at $t_{0}^{\prime}$, when the BH spin reaches the local maximum for the first time. Expanding the first term on the right side of Eq. (3.7) at $a_{*}=1$ and combining it with Eq. (3.8), Eq. (3.7) reduces to
$\dot{a}_{*}=$ $\frac{3\cdot 2^{1/3}}{\tau_{\text{acc}}}\left(1-a_{*}\right)^{2/3}-\frac{M_{\text{s,0}}}{24}M^{6}\mu^{8}\exp\left[\frac{\tau_{\text{acc}}}{192}(M^{8}-M^{8}_{0})\mu^{9}\right].$ (3.24)
@@
$\alpha_{0^{\prime}}=$ $\left(\frac{144}{\tau_{\mathrm{acc}}\mu}\right)^{{1}/{8}}\left[W_{0}\left(2^{{4}/{9}}3^{{2}/{3}}M_{\text{s,0}}^{-{4}/{3}}\tau_{\text{acc}}^{-{1}/{3}}\mu^{-{5}/{3}}e^{\frac{\alpha_{0}^{8}\tau_{\mathrm{acc}}\mu}{144}}\right)\right]^{{1}/{8}},$ (3.25)
@@
The time $t_{0}^{\prime}$ can then be calculated with Eq. (2.19), and the BH spin and condensate mass at $t_{0}^{\prime}$ can be derived by Eqs. (2.13) and (3.4), respectively. The subsequent evolution is the same as the scenario described in Sec. 3.2, with the subscript $0^{\prime\prime}$ in Eq. (3.11) replaced by $0^{\prime}$.
@@
Actually, Eq. (3.25) could be used to judge which one of the three groups the evolution should be attributed to. If the obtained $\alpha_{0^{\prime}}$ is larger than $\sqrt{6}k\alpha_{0}$, one should instead use the analysis in Sec. 3.2. On the other hand, if $\alpha_{0^{\prime}}$ is smaller than $\alpha_{0}$, the evolution does not follow the accretion trajectory at all, which case is discussed in Sec. 3.4 below.
3.4 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\gg{\tau}_{\mathrm{SR}}^{(011)}$
@@
This group includes three different scenarios ${\tau}_{\mathrm{SR}}^{(011)}\ll{\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{GW}}^{(011)}$, ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{GW}}^{(011)}$ and ${\tau}_{\mathrm{acc}}\gg{\tau}_{\mathrm{GW}}^{(011)}$, corresponding to ${f}_{\mathrm{Edd}}=10^{-9}$, $10^{-15}$ and $10^{-17}$ in Fig. 2, respectively. The curves with $\tau_{\text{acc}}\gg\tau_{\text{SR}}^{(011)}$ deviate from the $f_{\text{Edd}}=0$ case only from $t\sim\tau_{\text{acc}}$. The whole evolution can also be split into different phases. Specially, compared to the two cases analysed in Sections 3.2 and 3.3, the spin-up phase vanishes when $\tau_{\text{acc}}\gg\tau_{\text{SR}}^{(011)}$ and $a_{*0}>a_{*\text{c}}^{(011)}$.
@@
The BH spin drops from $t=0$ and reaches the minimum at time $t_{1}$. The BH mass $M_{1}$ and the condensate mass $M_{\text{s,1}}$ can be calculated in the same way as in Sec. 2.2.2. After $t_{1}$, the BH follows the Regge trajectory. The condensate mass may either increase or decrease, depending on the relative size of the superradiance rate and the GW emission rate at $t_{1}$. No matter which case, Eq. (3.18) can be used to estimate $\alpha_{2}$. If $\alpha_{2}>\alpha_{1}$, the condensate continues to grow after $t_{1}$ and $M_{\mathrm{s}}^{(011)}/M$ reaches its maximum value at $t_{2}$. Otherwise, the condensate shrinks since $t_{1}$. In the case of $\alpha_{2}>\alpha_{1}$, the time $t_{2}$ can be further calculated using Eq. (3.21), and the condensate mass at $t_{2}$ can be estimated with Eq. (3.16).
3.5 $a_{*0}<a_{*\text{c}}^{(011)}$
@@
When $a_{*0}$ is smaller than the critical value $a_{\text{c}}^{(011)}$, the superradiance condition is not satisfied initially. Nonetheless, the evolution of the BH-condensate system can still be split into the four phases explained above, as shown in Fig. 4. The angular momentum and energy flow from the condensate to the BH at first. Since the initial condensate mass is much smaller than the BH mass, this contribution is unimportant and the BH mass evolution is dominated by accretion. As a result, the growth of the BH mass can be estimated with Eq. (2.19). The BH spin also increases, following the relation in Eq. (2.13). Eventually, the BH spin is above the critical value and the superradiance is turned on. When superradiance balances accretion, the BH spin reaches its local maximum, which is the end of the spin-up phase. After the local maximum of the BH spin, the BH spin drops quickly to the Regge trajectory and then evolves along it afterward. The condensate mass first drops before the BH spin grows to its critical value $a_{*\text{c}}^{(011)}$. It can be estimated using Eq. (3.4) during this time interval. Since then, the system evolves as in the $a_{*0}>a_{\mathrm{*c}}^{(011)}$ cases, which are explained in previous subsections.
@@
In Fig. 4, we also compared the curves with $a_{*0}=0$ to those with $a_{*0}=0.7$. Because the condensate initial mass is small, the evolutions of BH mass are indistinguishable between these two cases, confirming our argument that the BH mass growth is dominated by accretion. In contrast, the BH spin and condensate mass evolution have different behavior with different values of $a_{*0}$. Especially, the spin-up and spin-down phases are largely affected by the initial BH spin. The corresponding estimates are similar to those in Sec. 3.3, with a revision to the calculation of $\alpha_{0^{\prime}}^{(011)}$. When $a_{*0}<a_{\mathrm{*c}}^{(011)}$, the spin-up phase can be divided into two segments, separated by the time $t_{\mathrm{int}}$ when the BH spin reaches $a_{*\text{c}}^{(011)}$. The BH mass at this time can be obtained by solving Eq. (2.10) together with Eq. (2.13). The mass coupling at this time, defined as $\alpha_{\text{int}}$, can be accomplished either numerically or analytically in powers of $\alpha_{0}$
$\frac{\alpha_{\text{int}}}{\alpha_{0}}=k+k\sum_{n=1}^{+\infty}c_{n}\left(\frac{k\alpha_{0}}{m}\right)^{n},$ (3.26)
where the coefficients $c_{n}$ can be obtained by solving Eq. (2.13) with Eq. (2.10) and Eq. (3.26) in orders of $\alpha_{0}$,
$c_{1}$ $=2^{5/2}3^{-3/2},$ (3.27a)
$c_{2}$ $=94/27,$ (3.27b)
$c_{3}$ $=\frac{2^{3/2}}{3^{7/2}}\left(172-9\frac{m^{2}}{\bar{n}^{2}}\right),$ (3.27c)
$c_{4}$ $=\frac{2}{729}\left(12451-1701\frac{m^{2}}{\bar{n}^{2}}\right).$ (3.27d)
@@
With the obtained $\alpha_{\text{int}}$, the condensate mass at $t_{\text{int}}$ can be estimated with Eq. (3.4). Beyond $t_{\text{int}}$, the formulas obtained above for the $a_{*0}>a_{\text{*c}}^{(011)}$ scenario can be applied. One could simply replace the $\alpha_{0}$, $a_{*0}$ and $M_{\text{s,0}}$ by $\alpha_{\text{int}}$, $a_{*\text{c}}^{(011)}$ and $M_{\text{s}}(t_{\text{int}})$, respectively.
@@
It is noteworthy that if the initial condensate mass is sufficiently large, the BH spin reaches ${a}_{\mathrm{*c}}^{(011)}$ at a faster rate with the additional contribution from the scalar field. Although this scenario can be easily adapted from Sec. 3.4, it is beyond the scope of this work. If modes with higher $l$ are considered, the BH spin may not reach $a_{*\text{c}}^{(011)}$. This scenario will be discussed in the next section.
@@
3.6 Evolution on the Regge plot
Figure 5: The evolution of the BH spin as a function of the BH mass. The initial parameters for panels (a) and (b) are identical to those used in Fig. 2 and Fig. 4, respectively. The $\{0,1,1\}$ Regge trajectory defined by Eq. (2.10) is represented as the upper boundary of the shaded region.
A classical BH free of charge has only two parameters: mass $M$ and spin $a_{*}$. The plot with $M$ as the horizontal axis and $a_{*}$ as the vertical axis is usually called Regge plot in the literature. Despite the complications in the evolution of the BH-condensate systems with different accretion rates, the evolution presents a rather simple and universal behavior on the Regge plot. In Fig. 5(a), we show the evolutions of the system with three typical values of $f_{\text{Edd}}$, corresponding to the three scenarios detailed in Sec. 3.2, Sec. 3.3 and Sec. 3.4, respectively. The initial BH spin and mass are $a_{*0}=0.7$ and $M_{0}=10^{3}M_{\odot}$, respectively. The initial condensate mass is $M_{\text{s},0}=10^{-5}M_{0}$ and the initial mass coupling is $\alpha_{0}=0.01$. In the evolution, only the $\{0,1,1\}$ mode is considered. The GW emission of the spinning condensate is also included. In Fig. 5(b), the initial BH spin is $a_{*0}=0$, corresponding to the scenario detailed in Sec. 3.5. In literature, the curves defined by Eq. (2.10) are commonly referred to as Regge trajectories, which correspond to the upper boundary of the shaded region in Fig. 5.
@@ 3.6 Evolution on the Regge plot
We then turn to the case with the fastest accretion, regardless of whether $a_{*0}>a_{*\mathrm{c}}^{(011)}$, corresponding to $f_{\text{Edd}}=1$ in Fig. 5. The accretion dominates the evolution of the BH and the superradiance can be safely ignored initially. The BH first follows the trajectory determined by Eq. (2.20), referred to as accretion trajectory below. The BH spin approaches $a_{*\text{lim}}$ and the BH mass grows almost exponentially. During this time, the condensate grows in mass and angular momentum. At some time, the condensate is dense enough such that the superradiance surpasses the accretion, leading to the fast drop to the Regge trajectory. The evolution follows the Regge trajectory afterward.
@@
From the Regge plot, it is clear that the entire evolution has a universal behavior. It consists of the parts following the two trajectories and the transition between them. They correspond to the spin-up, spin-down, and attractor phases in Sec. 3.2. The decaying phase could not be observed in the BH Regge plot. It is clear by looking at the evolution of the condensate mass.
@@
Due to this universal behavior, one only needs to identify the BH masse $M$ or spin $a_{*}$ at these critical points at which the system leaves the accretion trajectory and merges to the Regge trajectory. To determine these key points and the evolution between them, the following procedure can be applied. If $M_{0}\omega_{011}\geq 1/2$, superradiance does not occur and the system follows the accretion trajectory. If $M_{0}\omega_{011}<1/2$, $\alpha_{0^{\prime}}$ can be calculated using Eq. (3.25), as in Sec. 3 and Sec. 3.5. Based on the obtained $\alpha_{0^{\prime}}$, there are three possible cases:
@@
  • $\alpha_{0^{\prime}}>\sqrt{6}k\alpha_{0}$: This case corresponds to the scenario in Sec. 3.2 and the curves with $f_{\text{Edd}}=1$ in Fig. 5. The system initially follows the accretion trajectory approximately described by Eq. (2.13). Then it rapidly transits to the Regge trajectory in Eq. (2.10) at $\alpha_{1}$ given in Eq. (3.11). After that, the system follows the $\{0,1,1\}$ Regge trajectory until $M=M_{\mathrm{de}}^{(011)}$.
@@
  • $\alpha_{0}<\alpha_{0^{\prime}}<\sqrt{6}k\alpha_{0}$: This case corresponds to the scenario in Sec. 3.3 and the curves with $f_{\text{Edd}}=10^{-3}$ in Fig. 5. The system first follows the accretion trajectory and leaves this trajectory at $\alpha_{0^{\prime}}$. It then merges onto the Regge trajectories at $\alpha_{1}$ given in Eq. (3.11). Then it follows this trajectory until $M=M_{\mathrm{de}}^{(011)}$.
@@
  • $\alpha_{0^{\prime}}<\alpha_{0}$: This case corresponds to the scenario in Sec. 3.4 and the curves with $f_{\text{Edd}}=10^{-9}$ in Fig. 5. The system is dominated by superradiance initially and quickly merges to the Regge trajectories at $M=M_{1}$. Here $M_{1}$ is estimated using Eqs. (2.21) and (2.23b). The system then follows the $\{0,1,1\}$ Regge trajectory until $M=M_{\mathrm{de}}^{(011)}$.
@@
Finally, we turn to what is not reflected in the Regge plots. The BH mass grows almost exponentially before merging onto the Regge trajectory, from which one could obtain the corresponding time cost. This can then be applied to calculate the condensate mass and GW emission rate. When the BH evolves along the Regge trajectory, the Eqs. (3.12) to (3.23) can be directly applied for estimating the important quantities in this phase.
@@
With these equations, one could understand why the condensate mass evolution and GW emission flux are dramatically affected when $\tau_{\text{acc}}\lesssim\tau_{\text{SR}}$, which is shown in Fig. 2. Without accretion, the BH mass $M$ is roughly constant, resulting in an exponential increase of the condensate mass. The time duration of the GW is controlled by $\tau_{\text{GW}}$. Comparably, the condensate mass with accretion is a double exponential function and the termination of the GW emission is controlled by $\tau_{\text{de}}\ll\tau_{\text{GW}}$. The maximum condensate mass is also larger with accretion, which reaches as much as 0.15 in Fig. 3, while this number never exceeds 0.1 if accretion is absent. Consequently, the GW emission in Eq. (A.5) is stronger in magnitude but shorter in time.
4 Multi-mode evolution with accretion